OpenStudy (anonymous):

Does anyone know how to verify cos theta (tan squared theta + 1) = sec theta using the fundamental identities?

OpenStudy (bibby):

\[\sin^2(\theta) + \cos^2(\theta) = 1\]divide everything by cosine and you get: \[ \tan^2(\theta) + 1 = \sec^2(\theta)\]

OpenStudy (bibby):

Do you see how this applies to \[ \cos (\theta) (\tan ^2(\theta) + 1) = \sec (\theta)\]

OpenStudy (anonymous):

Yes thank you! :D

OpenStudy (bibby):

\[\huge :D\]

OpenStudy (anonymous):

I have so many questions I hope you don't mind?

OpenStudy (bibby):

I dunno if I can help on all of them but I'll try. You can always tag me

OpenStudy (anonymous):

Ok well do you know how to verify sin to the third x + sin x cos squared x = sin x using the fundamental identities?

OpenStudy (bibby):

Post another question please

OpenStudy (anonymous):

Simplify sin(pi/2-x)

OpenStudy (anonymous):

\[\sin(x-y)=\sin(x)\cos(y)-\cos(x)\sin(y)\]

OpenStudy (anonymous):

\[\sin(\pi/2-x)=\sin(\pi/2)\cos(x)-\cos(\pi/2)\sin(x)\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

\[\sin(\pi/2)=1 , \cos(\pi/2)=0\] \[\sin(\pi/2-x)=cos(x)\] Yw